OpenStudy (clare00):

For the following chemical reaction, what mass of hydrogen (in grams) will be produced from 4.23 mol of aluminum? 2Al + 3H2SO4 > 3H2 + Al2(SO4)3 in grams.

OpenStudy (kevin):

mass = mol x MR

OpenStudy (kevin):

mol of H2= 3/2* mol Al

OpenStudy (kevin):

mol of H2 = 3/2 * 4.23 = 6.345 mass of H2 =mol of H2 * Mr of H2